OpenStudy (anonymous):

solve 2cos^2(x)-tan(x)sin(x)=1-sec(x)

OpenStudy (anonymous):

Solve 2cos^2(x)-tan(x)sin(x)-1-sec(x) equal to what?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

im sorry it's suppose to be =1-sec(x)

OpenStudy (anonymous):

\[2\cos ^{2}x-\tan x \sin x=1-\sec x\]\[2\cos ^{2}x-(\sin x/\cos x)\sin x=1-1/\cos x\]\[2\cos ^{2}x-1/\cos x=1-1/\cos x \]\[2\cos ^{2}x=1\]\[\cos ^{2}x=1/2\]\[\cos x=\sqrt{1/2}\]\[\cos x=1\sqrt{2}\]\[x=pi/4\]

OpenStudy (anonymous):

Should be\[\cos x=\pm \sqrt{1/2}\]\[\cos x=\pm1/\sqrt{2}\]\[x=\pm pi/4\]